Mary left her estate worth $132,000 to three relatives. Her brother received 3 times as much as the cousin, and her daughter received 7 times as much as the cousin. How much did each relative receive?
Answer by Deina(147)

The cousin received C dollars.
Brother received 3C dollars and the
daughter received 7C dollars. So:

C + 3C + 7C = 132000

Now, combine like terms:
C + 3C = 4C; 4C + 7C = 11C
11C = 132,000

Now divide each side by 11:
C = 12,000

3C = 36,000
7C = 84,000 and bob's your uncle!